# Gaussian Splashing: Dynamic Fluid Synthesis with Gaussian Splatting

### v1

@@ 3.1. Position-based Dynamics
(2) $\Delta\bm{x}\leftarrow\bm{M}^{-1}\nabla\bm{C}^{\top}(\bm{x})\Delta\bm{\lambda}.$
@@
(3) $\Delta\bm{\lambda}_{j}\leftarrow\frac{-\Delta t^{2}C_{j}(\bm{x})-\bm{\alpha}_{j}}{\Delta t^{2}\nabla C_{j}\bm{M}^{-1}\nabla C_{j}^{\top}+\bm{\alpha}_{j}}.$
We note that such constraint-projection-based simulation naturally synergizes with 3DGS. It is also versatile and can deal with a wide range of physical problems such as fluid (Macklin and Müller, 2013; Xing et al., 2022), rigid bodies (Müller et al., 2020), or hyperelastic objects (Macklin and Muller, 2021).
3.2. 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) is a learning-based rasterization technique for 3D scene reconstruction and novel view synthesis. 3DGS encodes a radiance field using a set of Gaussian kernels $\mathcal{P}$ with trainable parameters $\left\{\bm{x}_{p},\sigma_{p},\bm{A}_{p},\bm{c}_{p}\right\}$ for $p\in\mathcal{P}$, where $\bm{x}_{p}$, $\sigma_{p}$, $\bm{A}_{p}$ and $\bm{c}_{p}$ represent the center, opacity, covariance matrix, and color function of each kernel. To generate a scene render, 3DGS projects these kernels onto the imaging plane according to the viewing matrix and blends their colors based on the opacity and depth. The final color of the $i$-th pixel is computed as:
(4) $\bm{c}_{i}=\sum_{k}G_{k}(i)\sigma_{k}\bm{c}_{k}(\bm{r}_{i})\prod_{j=1}^{k-1}\left(1-G_{j}(i)\sigma_{j}\right).$
Gaussian Shader (Jiang et al., 2023) further enhances 3DGS by incorporating additional trainable material parameters for kernel $p$ such as diffuse $\bm{d}_{p}$, specular $\bm{s}_{p}$, roughness $\rho_{p}$, and normal $\bm{n}_{p}$, along with a global environment map. It fuses more information into the kernel’s color:
(5) $\bm{c}_{p}(\bm{r}_{i})=\bm{d}_{p}+\bm{s}_{p}\odot L_{s}(\bm{r}_{i},\bm{n}_{p},\rho_{p}),$
4. Method
As shown in Fig. 2, the input of our system is a collection of images of a given 3D scene taken from different viewpoints. We separate foreground objects and the image background for all the inputs and extract the surface of masked objects. We apply an anisotropy loss to mitigate undesired splatting render to prevent over-stretching kernels when training 3DGS for the solid object. Doing so mitigates the rendering artifacts near the surface of solid models. On the other hand, fluids are made of spherical Gaussians. We track the normal for surface kernels to properly synthesize specular effects by augmenting them with an environment light map. Under the PBD framework, both fluids and solids can be animated in a unified way based on local constraint projection.
Figure 3. Different sampling strategies.  We compare the results of different sampling strategies: (left) fill the particle based on the density grid calculated using Gaussian kernels (Xie et al., 2023), and (right) uniformly sample within NeuS reconstruction. The point distribution generated by vanilla 3DGS is uneven, which hardly samples the legs or seat of the chair.
4.1. Scene Discretization
Gaussian kernels serve as the primary modality for discretizing geometry and dynamics, as well as for rendering the scene from novel camera poses. We exploit the state-of-the-art deep segmentation tools (Cheng et al., 2023; Kirillov et al., 2023) to separate foreground objects of interest out of the input images and obtain a dedicated object reconstruction. It is essential to have an accurate description of object boundaries in order to properly handle the object-object interaction or fluid-solid coupling. To this end, a surface mesh is built explicitly for each segmented foreground model using NeuS (Wang et al., 2021). NeuS extracts the zero-level set of a signed distance function corresponding to the foreground object. We find this method tends to yield a better-quality reconstruction compared with other possible alternatives.
We use the Poisson disk sampling (Bridson, 2007) to place Gaussian kernels inside the reconstructed model and regard each kernel as a mass particle for PBD simulation. In addition, we also assign kernels over the model’s surface for collision processing and fluid-solid interaction. It is noted that existing frameworks that incorporate physics with 3DGS (Xie et al., 2023) also need a spatial discretization over the model to numerically solve the dynamics equation. Those prior arts sample the model based on the learned density distribution. Doing so could result in sparsely sampled regions, especially for objects with thin parts, and potentially affect the simulation quality e.g., see Fig. 3.
4.2. Training
While the sampled kernels well represent the boundary of the objects and offer a simulation-ready discretization, their visual attributes or material parameters (e.g., diffuse, specular, roughness) have yet to be determined. Without them, high-quality rendering results are not possible. We leverage the differentiable 3DGS pipeline to optimize the appearance of every Gaussian kernel. All Gaussian kernels are first shaded with their corresponding material parameters (Eq. (5)) and are then splatted to a rendered image. Comparing it with the training view gives a loss back-propagated to update the corresponding parameters of Gaussian kernels. Specifically, the trainable parameters for each kernel $p$ are opacity $\sigma_{p}$, covariance $\bm{A}_{p}$ and material $\{\bm{d}_{p},s_{p},\rho_{p},\bm{n}_{p}\}$. During training we consider the kernel center $x_{p}$ as fixed to retain the discretization for the simulation. The loss is defined as:
(6) $\mathcal{L}=\mathcal{L}_{\rm color}+\lambda_{n}\mathcal{L}_{\rm normal}+\lambda_{a}\mathcal{L}_{\rm aniso}+\lambda_{v}\mathcal{L}_{\rm vol},$
where $\mathcal{L}_{\rm color}$ is the color loss between render and training image; $\mathcal{L}_{\rm normal}$ is normal consistency regularization adopted from Gaussian shader (Jiang et al., 2023). The regularization term $\mathcal{L}_{\rm aniso}$ is designed to prevent Gaussian kernels from becoming excessively elongated or compressed and potentially producing artifacts under large deformations. It is defined as:
(7) $\mathcal{L}_{\rm aniso}=\frac{1}{|\mathcal{P}|}\sum\limits_{p\in\mathcal{P}}{\max}\left\{\frac{\bm{S}_{p}^{1}}{\bm{S}_{p}^{2}}-a,0\right\},$
where $a$ is a ratio threshold and $\bm{S}_{p}=\left\{\bm{S}_{p}^{1},\bm{S}_{p}^{2},\bm{S}_{p}^{3}\right\}$ are the scalings of Gaussian kernels with $\bm{S}_{p}^{1}$ being the largest scaling and $\bm{S}_{p}^{3}$ being the smallest scaling. Note that as the shading normal is based on the minimal axis of Gaussian kernels, we do not constrain the minimal axis in the anisotropy loss. Otherwise, a spherical ellipsoid will result in normal ambiguity. We further encourage all Gaussian kernels to be relatively same-size by applying a regularization term:
(8) $\mathcal{L}_{\rm vol}=\frac{1}{|\mathcal{P}|}\sum\limits_{p\in\mathcal{P}}\left(\frac{\bm{S}_{p}^{1}\bm{S}_{p}^{2}\bm{S}_{p}^{3}}{\frac{1}{|\mathcal{P}|}\sum_{q\in\mathcal{P}}\bm{S}_{q}^{1}\bm{S}_{q}^{2}\bm{S}_{q}^{3}}-1\right)^{2}.$
We set $a=1.1$, $\lambda_{n}=0.2$, $\lambda_{a}=10$, and $\lambda_{v}=10$ in our experiments.
4.3. Position Based Fluids

### v2

@@ 3.1 Position-Based Dynamics
$\Delta\bm{x}=\bm{M}^{-1}\nabla\bm{C}^{\top}(\bm{x})\Delta\bm{\lambda}.$ (2)
We note that such constraint-projection-based simulation naturally synergizes with 3DGS. It is also versatile and can deal with a wide range of physical problems such as fluid [40, 70], rigid bodies [49], or hyperelastic objects [41].
3.2 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) is a learning-based rasterization technique for 3D scene reconstruction and novel view synthesis. 3DGS encodes a radiance field using a set of Gaussian kernels $\mathcal{P}$ with trainable parameters $\mathbf{x}_{p},\sigma_{p},\mathbf{A}_{p},\mathbf{c}_{p}$ for $p\in\mathcal{P}$, where $\bm{x}_{p}$, $\sigma_{p}$, $\bm{A}_{p}$ and $\bm{c}_{p}$ represent the center, opacity, covariance matrix, and color function of each kernel. To generate a scene render, 3DGS projects these kernels onto the imaging plane according to the viewing matrix and blends their colors based on the opacity and depth. The final color of the $i$-th pixel is computed as:
$\bm{c}_{i}=\sum_{k}G_{k}(i)\sigma_{k}\bm{c}_{k}(\bm{r}_{i})\prod_{j=1}^{k-1}\left(1-G_{j}(i)\sigma_{j}\right).$ (3)
GaussianShader [27] further enhances 3DGS by incorporating additional trainable material parameters for kernel $p$ such as diffuse $\bm{d}_{p}$, specular $\bm{s}_{p}$, roughness $\rho_{p}$, and normal $\bm{n}_{p}$, along with a global environment map. It fuses more information into the kernel’s color:
$\bm{c}_{p}(\bm{r}_{i})=\bm{d}_{p}+\bm{s}_{p}\odot L_{s}(\bm{r}_{i},\bm{n}_{p},\rho_{p}),$ (4)
4 Method
As shown in Fig. 2, the input of our system is a collection of images of a given 3D scene taken from different viewpoints. We separate foreground objects and the image background for all the inputs and extract the surface of masked objects. We apply an anisotropy loss to mitigate undesired splatting render to prevent over-stretching Gaussian kernels when training 3DGS for the solid object. Doing so mitigates the rendering artifacts near the surface of solid models. We decouple solid simulation and rendering by utilizing a separate set of sampled particles for simulation, and interpolating deformations of these particles onto trained Gaussian kernels during rendering. This approach ensures high-quality and robust results in both simulation and rendering. On the other hand, fluids use a unified set of Gaussian kernels (for rendering) or particles (for simulation). We track the fluids surface during simulation and use surface normal to properly synthesize specular effects by augmenting them with a decomposed environment map. Under the PBD framework, both fluids and solids are made of particles, and can be animated in a unified way based on local constraint projection.
4.1 Training
Our training process is generally similar to traditional Gaussian Splatting. However, due to our use of Physically-Based Rendering (PBR) modeling, the visual attributes or material parameters (e.g., diffuse, specular, roughness) of Gaussian kernels need to be determined. Without them, high-quality rendering results are not possible. Similar to [27], we leverage the differentiable 3DGS pipeline to optimize the appearance of every Gaussian kernel. All Gaussian kernels are first shaded with their corresponding material parameters (Eq. (4)) and are then splatted to a rendered image. Comparing it with the training view gives a loss back-propagated to update the corresponding parameters of Gaussian kernels. Specifically, the trainable parameters for each kernel $p$ are position $\bm{x}_{p}$, opacity $\sigma_{p}$, covariance $\bm{A}_{p}$ and material $\{\bm{d}_{p},s_{p},\rho_{p},\bm{n}_{p}\}$. The loss is defined as:
$\mathcal{L}=\mathcal{L}_{\rm color}+\lambda_{n}\mathcal{L}_{\rm normal}+\lambda_{a}\mathcal{L}_{\rm aniso},$ (5)
where $\mathcal{L}_{\rm color}$ is the color loss between render and training image; $\mathcal{L}_{\rm normal}$ is normal consistency regularization adopted from GaussianShader [27]. The anisotropy loss $\mathcal{L}_{\rm aniso}$ is designed to prevent Gaussian kernels from becoming excessively elongated or compressed and potentially producing artifacts under large deformations. It is defined as:
$\mathcal{L}_{\rm aniso}=\frac{1}{|\mathcal{P}|}\sum\limits_{p\in\mathcal{P}}{\max}\left\{\frac{\bm{S}_{p}^{1}}{\bm{S}_{p}^{2}}-a,0\right\},$ (6)
where $a$ is a ratio threshold and $\bm{S}_{p}=\left\{\bm{S}_{p}^{1},\bm{S}_{p}^{2},\bm{S}_{p}^{3}\right\}$ are the scalings of Gaussian kernels with $\bm{S}_{p}^{1}$ being the largest scaling and $\bm{S}_{p}^{3}$ being the smallest scaling. Note that as the shading normal is based on the minimal axis of Gaussian kernels, we do not constrain the minimal axis in the anisotropy loss. Otherwise, a spherical Gaussian kernel will result in normal ambiguity. We set $a=1.1$, $\lambda_{n}=0.2$, and $\lambda_{a}=10$ in our experiments.
4.2 Solid Simulation and Interpolation
In many dynamic Gaussian Splatting applications [69, 57], simulation and rendering are often coupled, utilizing the Gaussian kernels for both processes. This approach presents several issues. First, Gaussian kernels tend to distribute primarily on the surface of objects, necessitating the addition of internal kernels that, unless they contribute to effects like fracturing or breakage, are predominantly unrendered. Additionally, an uneven distribution of surface kernels can compromise the quality of the simulation, while a distribution that is too uniform can detrimentally affect rendering quality. Therefore, decoupling simulation and rendering is necessary to maintain quality in both aspects. In our pipeline, we leverage reconstructed 3DGS kernels for solid rendering and utilize a distinct set of particles for simulation.
To properly handle the object-object interaction or fluid-solid coupling, the particles are required to have an accurate description of object boundaries and interiors. To this end, we use the Poisson disk sampling [10] to place simulation particles within the surface mesh, which is explicitly reconstructed for segmented foreground model using NeuS [65]. NeuS extracts the zero-level set of a signed distance function corresponding to the foreground object. It is important to note that existing frameworks that incorporate physics with 3DGS [69] also require spatial discretization over the model to numerically solve the dynamics equation. Those prior approaches sample the model based on the trained 3DGS kernels. This method can result in sparsely sampled regions, especially for objects with thin parts, potentially affecting simulation quality (see Fig. 11 in supplementary material).
Once the simulation particles are sampled and in place, we perform PBD on them to animate frames of dynamic motion. Each frame, deformation gradients and displacements are interpolated from the simulation particles to the trained Gaussian kernels using generalized moving least squares (GMLS) [43], which produces smooth and robust results.
4.3 Position-Based Fluids
We employ the Position-Based Fluids (PBF) [40] as our Lagrangian fluid synthesizer. To enforce the fluid incompressibility, PBF imposes a density constraint $C^{\rho}_{i}$ on each particle, maintaining the integrated density $\rho_{i}$ computed by the SPH kernel as:
$C_{i}^{\rho}=\frac{\rho_{i}}{\rho_{0}}-1=\sum_{j}\frac{m_{j}}{\rho_{0}}W(\bm{p}_{i}-\bm{p}_{j})-1,$ (7)
where $m_{j}$ is the mass of particle $j$. $\bm{p}_{i}$ is the position of particle $i$, and $W$ is the SPH kernel function. Each constraint can be straightforwardly parallelized using GPUs.
GSP incorporates a position-based tension model [70] to more accurately simulate fluid surface dynamics. Surface detection for a particle is performed through occlusion estimation, where each particle is enclosed by a spherical screen. Neighboring particles project onto this screen, and a particle is classified as on the surface if the cumulative projection area is below a specified threshold, reflecting its partial exposure. Given the natural tendency of tensions to reduce surface area, PBF enforces an area constraint on each surface particle to minimize local surface area. This process begins with the calculation of the normal $\bm{n}_{i}$ for surface particle $i$ as:
$\bm{n}_{i}=\text{normalize}(-\nabla_{\bm{p}_{i}}C_{i}^{\rho}),$ (8)
$C_{i}^{A}=\sum_{t\in T(i)}\frac{1}{2}\left\|(\bm{p}_{t^{2}}-\bm{p}_{t^{1}})\times(\bm{p}_{t^{3}}-\bm{p}_{t^{1}})\right\|$ (9)
$C_{ij}^{D}={\min}\left\{0,\left\|\bm{p}_{i}-\bm{p}_{j}\right\|-d_{0}\right\},$ (10)
where $d_{0}$ is the distance threshold. The original version was parallelized on CPUs; we have enhanced it for GPU parallelization. Calculations for $\bm{n}_{i}$, $C_{i}^{A}$ and $C_{ij}^{D}$ are efficiently parallelizable on GPUs. Considering that the number of surface particles surrounding each particle is typically low, local triangulation for each particle is conducted independently on separate threads.
4.4 Rendering
$\bm{A}_{p}^{t}=\bm{F}_{p}\bm{A}_{p}\bm{F}_{p}^{\top},\;\text{and}\;\bm{n}_{p}^{t}=\frac{\bm{F}_{p}^{-\top}\bm{n}_{p}}{\left\|\bm{F}_{p}^{-\top}\bm{n}_{p}\right\|}.$ (11)
We then shade the deformed Gaussian kernels $\{\bm{x}^{t}_{p},\sigma_{p},\bm{A}_{p}^{t},\bm{n}_{p}^{t}\}$ with material $\{\bm{d}_{p},s_{p},\rho_{p}\}$, i.e., Eq. (4) and splat them into an image $\bm{c}^{\rm bg}$. As shadows are important to visual outcomes in dynamic scenes, we further re-engineer nearly-soft shadows [14] into our system to enhance the realism.
Figure 3: GSP rendering.  GSP synthesizes high-quality images corresponding to dynamically interacting fluids and solids. (a) The final rendered image combining rendered solids, fluids, and foams. (b) The rendering result of solids. (c) The rendering result of fluids. (d) The fluid thickness by additive splatting, where the darker color indicates the higher thickness. (e) The normal of fluids. (f) The intensity of foam particles. The insets in (b) and (e) represent solid and fluid Gaussian kernels, respectively.
The dynamic particle-based fluids are rendered with ellipsoids splatting [40], which is inherently compatible with the existing 3DGS pipeline. We begin by generating spherical fluid Gaussian kernels at each fluid particle. The initial covariance $\bm{A}_{p}$ of each kernel is determined by the particle radius. The normal $\bm{n}_{p}$ adopts the surface normal of the nearest surface fluid particle from PBF simulation. We proceed to set up the appearance material for each fluid Gaussian kernel, which requires careful engineering due to the fluid’s strong reflection and refraction effects. We employ the current PBR workflow to model the reflection, which adopts the same formula of Eq. (4). We set a specular material ($s_{p}=1$, $\rho_{p}=0.05$ in our experiments) for all fluid kernels to imitate reflective behavior. The refraction needs careful treatment, which we model into a thickness-dependent diffuse color $\bm{d}_{p}$. As the fluid thickness increases, the absorption of light within the fluid intensifies, resulting in reduced visibility for objects behind. Conversely, when there is less fluid present, it exhibits a more transparent appearance. The fluid thickness $\tau$ comes from the modified splatting pipeline, with the alpha blending replaced by additive blending. The refraction color $\bm{d}_{p}$ is then represented by Beer’s Law [62]:
$\bm{d}_{p}=e^{-k\tau_{p}}\bm{c}_{p}^{\rm bg}.$ (12)
Here, the absorption coefficient $k$ is defined differently for each color channel, $\tau_{p}$,$\bm{c}_{p}^{\rm bg}$ is the fluid thickness and background back-projected to each Gaussian kernel respectively. Note that for background back-projection, a distortion $\beta\bm{n}_{p}$ is added to mimic the change of light path due to refraction. Opacity $\sigma_{p}$ is set to $1$ as most of transmission and refraction has already been modeled into $\bm{d}_{p}$. We finally shade all fluid Gaussian kernels $\{\bm{x}^{t},\sigma_{p},\bm{A}_{p}^{t},\bm{n}_{p}^{t}\,\bm{d}_{p},s_{p},\rho_{p}\}$ and splat them to the fluid rendering result $\bm{c}^{\rm fluid}$ with Gaussian Splatting. To enhance the realism of fluids, foam particles are synthesized [23] and rendered [1]. The 3DGS pipeline is re-engineered to incorporate additive splatting with distinct kernels for foam, bubble, and spray particles. The final rendering result is achieved by combining the $\bm{c}^{\rm bg}$ and $\bm{c}^{\rm fluid}$, as shown in Fig. 3. Please refer to the supplementary material for more details on the rendering part.
4.5 Inpainting
Displaced object exposes unseen areas that were originally covered to the camera. Since they are not present in the input image, 3DGS is unable to recover the color and texture information of these areas, leading to black smudges and dirty textures in the result. GSP remedies this issue using an inpainting trick. First, we remove all the Gaussian kernels of the object that may be displaced. We then use LaMa [61] to inpaint the rendered results, noting the new colors of pixels originally in spots and assigning them to the first Gaussian kernel encountered by the rays emitted from these pixels. We average the recorded colors on all noted Gaussian kernels for their diffuse color, set their opacity to $1$, and their specular color to $\bm{0}$ to prevent highlights.
5 Experiments
We implemented Gaussian Splashing pipeline using Python, C++ and CUDA on a desktop PC equipped with a 12-core Intel i7-12700F CPU and an NVIDIA RTX 3090 GPU. Specifically, for the rendering part, we ported the published implementation of GaussianShader [27] and integrated our fluid rendering using PyTorch [24]. PBD/PBF engine was implemented with CUDA, and we also group independent constraints to efficiently parallelize the constraint projections on the GPU. Please refer to the supplementary material for more details of the implementation.
5.1 Ablation
Figure 4: Anistropy regularization.  Anistropy regularization effectively maintains rendering quality under large deformations. Without the regularization term, 3DGS tends to generate fuzzy and spiky artifacts, especially near the surface of the model (left). When the regularization is applied, image quality is greatly improved with correct specular effects.
Anisotropy Regularization
3DGS is originally designed for view synthesis. 3DGS obtained from a static scene produces low-quality renders when Gaussian kernels undergo large rotational deformations. The anisotropy regularization is effective against this limitation as shown Fig. 4. Detailed statistics regarding the experiment settings and timings are reported in Tab. 1. Most experiments can also be found in the supplementary video.
PBR Material
To show the importance of specular highlights in fluid rendering, we show a side-by-side comparison in a 3DGS scene. As shown in Fig. 5, we compare fluids rendered using purely diffuse materials with those that incorporated specular reflections. The fluids without specular reflection appear almost smoke-like, while the inclusion of specular term significantly enhances the realism of the fluids. It should be noted that incorporating PBR in 3DGS is key to improving the render quality of largely deformed fluids. This is in contrast to previous work [69], which baked lighting into spherical harmonics and failed to produce realistic rendering when fluids underwent significant motion.
Figure 5: Ablation study of specular.  We demonstrate the impact of specular highlights on the quality of rendering. On the left is a fluid rendered with diffuse color only. On the right, surface reflective specular are added, which exhibits a more realistic and dynamic fluid.
Shadow Map
In Fig. 6, we demonstrate that GSP, with the addition of nearly soft shadows, significantly improves visual realism compared to vanilla 3DGS. Without these shadows, objects appear like flat layers pasted onto the background, lacking a sense of depth.
Figure 6: Shadow map. GSP incorporates dynamic shadows into the rendering pipeline to enhance visual realism. We re-engineer variance shadow mapping [14] within the existing 3DGS pipeline to produce nearly-soft shadows.
Inpainting
We conduct an ablation experiment on inpainting as shown in Fig. 7. In this indoor scene, displacing cup and dog with vanilla 3DGS results in the occurance of black smudges and dirty textures, as the hidden area by the object could not be reconstructed properly due to missing information in the input images. GSP addresses this by using LaMa [61] for inpainting.
Figure 7: 3DGS inpainting.  In this indoor scene, both the paper cup and the stuffed toy dog are segmented from the input image (left). 3DGS leaves empty spots and dirty textures blended from irrelevant kernels, as highlighted in the middle figure. Applying the inpainting with generative AI [61] ameliorates this issue (right).
5.2 Evaluation
We evaluate GSP through a diverse set of experiments, covering the dynamics of deformable bodies, rigid bodies, and fluids, as well as two-way coupling between solids and fluids. For additional frames and more detailed results, please refer to the supplementary material. All experiments are also available in the supplementary video. A preliminary test is depicted in Fig. 12, where a soft chair from the NeRF synthetic datasets [44] falls into a pool, demonstrating the two-way coupling between deformable solids and fluids. The chair deforms, floats due to buoyancy, and generates fluid ripples. Fig. 13 illustrates another dynamic fluid scene featuring a coastal cliff and waves. The waves continuously push towards the shore from a distance, and upon colliding with the cliff, they splash out foam and spray. This accurately models the complex interactions between the fluid and the solid cliff face. Fig. 14 showcases another fluid-solid interaction test. The garden scene is sourced from the Mip-NeRF 360 [5] dataset. The foreground objects consist of a fixed table and a potted plant. We pour water into the garden slowly. The water rises up and eventually sinks the table, and sweeps the plant. In Fig. 8 and Fig. 15, a lego bulldozer is surfing on the splashing waves. Through the two-way coupling, the baseplate and the bucket deform and vibrate under the impact of the waves.
GSP has a semantic segmentation module. Therefore, the user is able to freely manipulate the models in the scene. Furthermore, since everything is represented as Gaussian particles, GSP allows the user to transform the state of the model. An example is shown in Fig. 16, the scene includes a round white table, on which a paper cup and a stuffed toy dog are placed. Water is poured into the cup. The state of the toy dog and the cup are changed to water, and splashes on the desk. As mentioned, we use LaMA to inpaint the table texture so that the user does not observe rendering artifacts when the liquefied cup and toy dog splash away.
Table 1: Time performance.  We present detailed time statistics for the experiments reported in the paper. All time-related evaluations are expressed in seconds. From left to right, (1) # Kernels indicates the number of Gaussian kernels, while # Solids, # Fluids, and # BG denote the foreground solid, the fluid, and the background, respectively. In some experiments (e.g., Cup & Dog and Headset), the number of fluid particles varies over time. We report the maximum number of fluid particles. (2) Sim. Time provides time statistics for simulations, with Overall representing the total simulation time per time step and Tension indicating the time taken for each surface tension constraint projection. (3) Render Time details the time statistics for rendering, with Solids, Fluids, and BG denoting the time spent on rendering the solids, fluids, and background, respectively.
Scene (Fig.) # Kernels Sim. Time (s) Render Time (×$\times 10^{-2}$ s)
# Solids # Fluids # BG Overall Tension Solids Fluids BG
Chair (Fig. 12) 315K 300K 0 5.4 1.1 3.9 5.9 0
Waves (Fig. 13) 420K 817K 0 8.1 3.1 8.2 11.6 0
Garden (Fig. 14) 450K 614K 2.27M 7.3 1.6 6.9 10.3 2.3
Lego (Fig. 8&15) 330K 280K 290K 3.8 1.0 2.9 4.6 1.9
Cup & dog (Fig. 16) 156K 160K 310K 2.1 0.6 2.4 4.1 1.9
Headset (Fig. 9&17 ) 357K 64K 2.22M 1.9 1.8 1.5 3.8 2.7
Can (Fig. 18) 390K 254K 1.19M 1.6 0.8 2.5 4.7 2.0
Astronaut (Fig. 19) 0 145K 0 0.8 0.6 0 4.0 0
Ficus (Fig. 20) 204K 0 0 2.3 0 2.6 0 0
Bulldozers (Fig. 21) 6.67M 0 350K 4.5 0 24.6 0 2.1
Figure 8: Splashing LEGO.  Through the two-way coupling dynamics, the LEGO bulldozer is animated to surf on the splashing waves.
Figure 9: Headset droplets. Water flows from a headset hanging above an office desk, resembling a faucet. Due to surface tension, the water forms droplets, sliding down the computer screen and splashing onto the desk, creating a puddle.
Our fluid simulator can also capture the surface tension within the PBD framework. This enables realistic low-volume fluid-solid interaction. As shown in Fig. 9 and Fig. 17, water flows from a headset hanging above an office desk, resembling a faucet. Due to surface tension, the water forms droplets as it falls, sliding down the computer screen and splashing onto the desk, creating a puddle. In Fig. 18, droplets of water continuously drip onto the top of the can until they reaches the capacity and overflow. The surface tension of the liquid causes the droplets to gradually aggregate at the surface. As more droplets fall onto the liquid surface, they rise above the edge of the can. Eventually, the accumulated water exceeds the limit of surface tension and spills over. Fig. 19 showcases another interesting use of GSP, where the user applies Trisolarans’ black magic on an astronaut. The astronaut is strucked by the magic and is transformed to water. It finally collapses into a water ball in a zero-gravity space due to the presence of surface tension.
GSP is a versatile system and supports the manipulation of both rigid objects and deformable bodies. As shown in Fig. 20, a deformable ficus is waving at you. Due to the external force applied to it, the plant undergoes continuous deformation. In Fig. 21, bulldozers are piled and dropped in a bowl. They collide and contact with each other and eventually scattered within the box.
6 Conclusion
GSP is a novel pipeline combining versatile position-based dynamics with 3DGS. The principle design philosophy of Gaussian Splashing is to harness the consistency of volume particle-based discretization to enable integrated processing of various 3D graphics and vision tasks, such as 3D reconstruction, deformable simulation, fluid dynamics, rigid contacts, and rendering from novel viewpoints. While the concept is straightforward, building Gaussian Splashing involves significant research and engineering efforts. The presence of fluid complicates the 3DGS processing due to the specular highlights at the fluid surface. Fluid-solid coupling resorts to accurate surface information; Large deformation on the solid object generates defective rendering; Displaced models also leave empty regions that the input images fail to capture. We overcome those difficulties by systematically integrating and adapting a collection of state-of-the-art technologies into the framework. As a result, Gaussian Splashing enables realistic view synthesis not only under novel camera poses but also with novel physically-based dynamics for various deformable, rigid, and fluid objects, or even novel object state transform. It should be noted that incorporating physically-based fluid dynamics in NeRF/3DGS has not been explored previously. The primary contribution of this work is to showcase the feasibility of building a unified framework for integrated physics and learning-based 3D reconstruction. Gaussian Splashing still has many limitations. For instance, PBD is known to be less physically accurate. It may be worth generalizing PBD with other meshless simulation methods. The fluid rendering in Gaussian Splashing in its current form is far from perfect – ellipsoid splatting is an ideal candidate for position-based fluid but does not physically handle real world light transport, e.g. refraction.
Supplementary Material
7 Simulation Details
7.1 Position-Based Dynamics
PBD/XPBD treats a dynamic system as a set of $N$ vertices, i.e., $\bm{x}=[\bm{x_{0}},\bm{x_{1}},...,\bm{x_{N}}]^{\top}$ and $M$ constraints, i.e., $\bm{C}(\bm{x})=[C_{1}(\bm{x}),C_{2}(\bm{x}),...,C_{M}(\bm{x})]^{\top}$. Here, $\bm{x}$ represents the position of vertices and $\bm{C}(\bm{x})$ represents the set of constraints. Specifically, the total system potential $U$ is defined as a quadratic form of all the constraints such that $U=\frac{1}{2}\bm{C}^{\top}(\bm{x})\bm{\alpha}^{-1}\bm{C}(\bm{x})$. Here, $\bm{\alpha}$ is the compliance matrix, i.e., the inverse of the constraint stiffness. For example, if there are only two vertices and they form a mass-spring system, constraint and compliance matrix could be written as $\bm{C}(\bm{x})=[\|\bm{x_{0}}-\bm{x_{1}}\|-d_{0}]^{\top}$ and $\bm{\alpha}=[k]$, where $d_{0}$ is the rest length of the spring and $k$ is the stiffness of the spring.
Motion at each time step can be solved by minimizing the system energy. However, PBD/XPBD offers an easy and efficient simulation modality, converting the variational optimization to the so-called constraint projections.
XBPD estimates an update of constraint force (i.e., the multiplier) $\Delta\bm{\lambda}$ by solving:
$\left[\Delta t^{2}\nabla\bm{C}(\bm{x})\bm{M}^{-1}\nabla\bm{C}^{\top}(\bm{x})+\bm{\alpha}\right]\Delta\bm{\lambda}=-\Delta t^{2}\bm{C}(\bm{x})-\bm{\alpha}\bm{\lambda},$ (13)
$\Delta\bm{x}=\bm{M}^{-1}\nabla\bm{C}^{\top}(\bm{x})\Delta\bm{\lambda}.$ (14)
@@ 7.1 Position-Based Dynamics
$\Delta\bm{\lambda}_{j}\leftarrow\frac{-\Delta t^{2}C_{j}(\bm{x})-\bm{\alpha}_{j}}{\Delta t^{2}\nabla C_{j}\bm{M}^{-1}\nabla C_{j}^{\top}+\bm{\alpha}_{j}}.$ (15)
A typical XPBD simulation loop is shown in Algorithm 1.
Algorithm 1 XPBD simulation loop for time step $n+1$
1:predict position $\tilde{\bm{x}}\Leftarrow\bm{x}^{n}+\Delta t\bm{v}^{n}+\Delta t^{2}\bm{M}^{-1}\bm{f}_{ext}(\bm{x})^{n}$
2:
3:initialize solve $\bm{x}_{0}\Leftarrow{\bm{x}}$
4:initialize multipliers $\bm{\lambda}_{0}\Leftarrow{\bm{0}}$
5:while $i<{\rm solverIterations}$ do
6:       for all constraints do
7:             compute $\Delta\lambda$ using Eq. 15
8:             compute $\Delta\bm{x}$ using Eq. 14
9:             update $\lambda_{i+1}\Leftarrow\lambda_{i}+\Delta\lambda$
10:             update $\bm{x}_{i+1}\Leftarrow\bm{x}_{i}+\Delta\bm{x}$
11:       end for
12:end while
13:update positions $\bm{x}^{n+1}\Leftarrow\bm{x}_{i}$
14:update velocities $\bm{v}^{n+1}\Leftarrow\frac{1}{\Delta t}(\bm{x}^{n+1}-\bm{x}^{n})$
7.2 Position-Based Fluids